A metal worker has several 1-kilogram bars of a metal alloy that contain 23% copper and several 1-kilogram bars that contain 79% copper. How many bars of each type of alloy should be melted and combined to create 48 kilograms of a 44% copper alloy? Set up and solve a system of equations to find the answer.

Answer by stanbon(75887)

A metal worker has several 1-kilogram bars of a metal alloy that contain 23% copper and several 1-kilogram bars that contain 79% copper. How many bars of each type of alloy should be melted and combined to create 48 kilograms of a 44% copper alloy? Set up and solve a system of equations to find the answer.
-----------
Quantity Eq:: t + s = 48 kg
Copper Eq:: 23t+79s = 44*48
---------------------------------
Modify for elimination::
23t + 23s = 23*48
23t + 79s = 44*48
---------------
Subtract and solve for "s"::
56s = 21*48
----
s = 18 (# of 79% bars needed)
t = 48-18 = 30 (# of 23% bars needed)
-----------------
Cheers,
Stan H.
----------------